At what speed does a 1300 kg compact car have the same kinetic energy as a 1.8×104 kg truck going 24 km/hr ?
Answer #1

m1v12/2=m2v22/2 <--> v_{1}=v_{2}\sqrt{\frac{m_{2}}{m_{1}}}

Num: v1~89.305 km/h